# Zhiyuan Liu3, Jinan Xu1, Yuhua Li2, Ruixuan Li2, Maosong Sun3

### v2

@@ Abstract
The recent advancement of Multimodal Large Language Models (MLLMs) has significantly improved their fine-grained perception of single images and general comprehension across multiple images. However, existing MLLMs still face challenges in achieving precise grounding in complex multi-image scenarios. To address this, we first explore a Chain-of-Thought (CoT) framework that integrates single-image grounding with multi-image comprehension. While partially effective, it remains unstable and struggles to capture abstract visual information due to its non-end-to-end nature. Therefore, we introduce Migician, the first multi-image grounding model capable of performing free-form and accurate grounding across multiple images. To support this, we present the MGrounding-630k dataset, which comprises data for several multi-image grounding tasks derived from existing datasets, along with newly generated free-form grounding instruction-following data. Furthermore, we propose MIG-Bench, a comprehensive benchmark specifically designed for evaluating multi-image grounding capabilities. Experimental results demonstrate that our model achieves significantly superior multi-image grounding capabilities, outperforming the best existing MLLMs by 21.61% and even surpassing much larger 70B models. Our code, model, dataset, and benchmark are fully open-sourced at https://migician-vg.github.io/.
@@ 1 Introduction
As a result, a question naturally arises: Can we integrated the single-image grounding and multi-image understanding capabilities of existing MLLMs to tackle the MIG task? In this work, we propose a Chain-of-Thought (CoT) framework that first leverages multi-image understanding to generate a textual referring query, and then utilizes it for localization through single-image grounding. This approach is proven highly effective for MIG tasks, particularly in simple scenarios where textual descriptions are sufficiently distinctive, revealing the potential of MLLMs in handling such tasks.
However, the proposed CoT framework struggles with describing abstract visual semantics in multi-image scenarios, and the two-step process results in a doubling of the inference time. To address this, we further propose Migician, a competitive MLLM capable of free-form and accurate grounding across multiple images, which is an end-to-end solution for MIG. To progressively establish flexible grounding capabilities, we employ a two-stage training procedure based on our proposed large-scale MIG dataset (MGrounding-630k). First, the grounding ability of Migician is enhanced through a combination of data of MIG tasks and general tasks. Then, Migician is further refined using high-quality free-form MIG instruction data. In addition, to evaluate the challenges of the free-form MIG scenario, we construct a comprehensive multi-image grounding benchmark, MIG-bench, comprising a total of 10 different tasks, 5.9k diverse images and more than 4.2k test instances. We observe a significant gap between the performance of existing mainstream MLLMs and human performance on the MIG-bench. In contrast, Migician can effectively alleviate this gap and improve the performance of free-form MIG.
@@ 4 Methods
In this section, we delve into the methods for enabling free-form multi-image grounding capabilities in MLLMs. Since free-form MIG requires the ability to perform visual grounding while simultaneously understanding multiple images, we begin by investigating a Chain-of-Thought (CoT) framework to combine these two capabilities within existing MLLMs to tackle this task. Furthermore, we develop an end-to-end MIG model, Migician, through instruction tuning to overcome the limitations of the CoT framework and achieve enhanced MIG performance.
@@ 4.1 A Chain-of-Thought Framework
Although some existing MLLMs such as Qwen2-VL-7B [38] demonstrate strong multi-image understanding and single-image grounding capabilities, we find that directly prompting them to perform MIG tasks often leads to significant performance degradation as illustrated in Figure 3(a). To better explore the potential of existing models for MIG tasks, we design a Chain-of-Thought (CoT) framework that enables the model to effectively leverage and combine its exitsing abilities during the MIG execution.
Specifically, we decompose the MIG task into two subtasks as illustrated in Figure 3(b). The model is first prompted to engage in a "reasoning process" by performing multi-image understanding based on the input images and the given prompt, generating a textual referring expression that describes the target object. Next, the model performs the visual grounding task, using the referring expression from the previous step to locate the objects in corresponding images. This framework leads to a notable performance improvement on MIG tasks, indicating that existing MLLMs possess the underlying capabilities required for such tasks but need an effective method to elicit them.
However, the CoT framework suffers from several inherent limitations. On one hand, the multi-step process introduces error propagation issues [44] and impacts reasoning efficiency. On the other hand, many scenarios require grounding through abstract visual semantics across multi-image contexts (as shown in Figure 3(c)), making the use of an intermediate textual referring expression impractical. This highlights the need for an end-to-end model capable of directly performing the MIG task.
More failure patterns of the CoT framework are illustrated in Figure 5, categorized into perceptual and reasoning flaws. For the former, the framework falls short when multiple images are organized in a manner where only integrating all their visual information could tackle MIG (i.e. finding the location of missing people in the second image), or when the textual content could not sufficiently represent the visual information. Regarding reasoning errors, inaccuracies can arise at various stages of the reasoning process, undermining the framework’s overall accuracy and effectiveness.
@@ 4.2 Data Construction
By analyzing the tasks and annotation types of existing datasets, we identify multiple multi-image grounding (MIG) tasks whose data could be derived through transformation of the existing. Specifically, we collect and organize data from existing sources, combining or automatically synthesizing single-image annotations to create datasets for 6 types of MIG tasks. Each task contains over 70k examples, resulting in a total of 530k training samples. The details of these task data are in Appendix C.1.
@@
Synthesizing Free-form MIG Data.
The data obtained through the aforementioned methods still do not fully meet the requirements for free-form MIG. To acquire MIG data with richer and more diverse formats, which would enhancing the model’s instruction-following and flexible grounding capabilities, we design a MIG data synthesis pipeline. This pipeline uses the Objects365 [35] images with object annotations, select multiple images as a group, and generate high-quality instructions for multi-image grounding. Specifically, we first employ Qwen2-VL-72B [38] to generate captions of each individual image and then perform error filtering and refinement on the annotated bounding boxes. Next, we prompt Qwen2.5-72B [43] to automatically generate high-quality, free-form MIG question-answering pairs by integrating information from multiple images. To optimize the selection of appropriate image groups, we adopt different image grouping methods, including random selection, selection of images with common objects, and grouping images based on CLIP similarity to select semantically similar images for each. Using these methods, we generate a total of 100k Free-Form MIG data. For more detailed information please refer to Appendix C.2.
4.3 Instruction Tuning for MIG
Two-Stage Training.
To effectively equip the model with free-form MIG capabilities, we propose a two-stage training approach. In the first stage, the model learns to perform multi-image grounding by training on the six representative MIG tasks of MGrounding-630k, acquiring the ability to simultaneously comprehend multiple images and execute visual grounding. In the second stage, the model is further fine-tuned on free-form MIG instruction data in MGrounding-630k, enabling it to adapt to more flexible and diverse instruction types and transfer the MIG skills learned in the first stage to a broader range of scenarios. To prevent the model from forgetting its existing capabilities during training, we also incorporate single-image understanding, multi-image understanding, and single-image grounding data into each training stage. More details are in Appendix D.
@@ 4.3 Instruction Tuning for MIG
After the second stage of fine-tuning, we observe a trade-off between model performance and flexibility: while the model adapts to the free-form MIG instructions, there is a performance drop in common multi-image grounding tasks. To better balance these two aspects, we adopt the model merging technique [14], averaging the model weights obtained from stage-2 with different training settings as the final weights. We find this approach mitigates the performance loss in common MIG tasks while preserving the ability to follow free-form MIG instructions effectively.
@@
V* Bench Attribute Spatial Overall
Human Level 98.26 100.00 98.95
Random Guess 26.73 50.00 35.99
Tool-using Pipeline
MM-React 34.78 51.31 41.36
Visprog 31.30 56.57 41.36
SEAL 74.78 76.31 75.39
End-to-end MLLMs
InternVL2-8B 29.56 56.57 43.07
Gemini Pro 40.86 59.21 48.16
LLaVA-1.5 43.47 56.57 48.68
Minicpm2.6 40.86 64.47 52.67
GPT-4V 51.30 60.53 54.97
Migicianzero_shot 59.16 60.53 59.85
Migicianslice 77.49 67.11 72.30
Table 3: On V* Bench, Migician generalizes well to the hyper-resolution single image in a zero-shot manner.
5 MIG-Bench
To thoroughly assess the multi-image grounding abilities of current MLLMs, we have meticulously curated MIG-Bench. This benchmark consists of 5.9k images and 4.3k testing instances, covering 10 tasks. The distribution of these tasks is illustrated in Figure 4. The benchmark tasks are divided into two categories: spontaneous and referential multi-image grounding. Spontaneous grounding tasks require the model to recognize and ground differences or common objects across images, with a total of 1.4k testing instances. Referential grounding tasks, on the other hand, require the model to utilize different forms of reference queries (i.e., textual, visual, or multimodal) to locate the target objects, comprising 6 distinctive tasks and 2.9k testing instances. The details of these tasks are provided in Figure 2 and Appendix A.
To ensure diversity, the images are sourced from a variety of sources, existing datasets, web images and manually captured photos. For existing datasets, we use examples that exhibits significant movement from GOT-10k_val [13] for the Object Tracking task, and manually modify the images from Objects365 [35] for Common Object task. For Multi-view Grounding, we collect 288 examples spanning both indoor and outdoor scenes from Ego4D [11]. The Static Difference task is sourced from the MagicBrush_dev set [50]. We also manually capture 97 image pairs with view differences in real-world settings and collect, on average, over 100 image pairs from Google Search for tasks related to Visual Referring, Reasoning, and Referring Grounding tasks.
For a quantitative and objective evaluation of different MLLMs, we ensure that each testing instance contains only one clear target region, eliminating any potential ambiguity. Our MIG-Bench offers a comprehensive evaluation across various real-world scenarios and domains. We believe this benchmark will provide valuable insights into the challenges of MIG and inspire further research in related areas.
6 Experiments
6.1 Implementation Details
Migician undergoes development based on the Qwen2-VL-7B [38] foundation model with a global batch size of 48, a total of 25,000 steps for the two-stage training procedure, and a learning rate of 5e-6, using 8×A100-80G GPUs. For the evaluation in our proposed MIG-Bench, we use the conventional metric $\text{Acc}_{0.5}$ in referring expression comprehension [19]. This metric measures the accuracy of object localization, defining a prediction as correct if the Intersection over Union (IoU) with the ground truth bounding box is greater than 0.5.
@@ 6.1 Implementation Details
Setting Multi-image General Benchmarks Multi-image Grounding
MuirBench BLINK MIBench Mantis MMIU MIG
@@
Table 6: The ablation study about the contribution of different data subsets.
6.2 Results on MIG-Bench
As shown in Table 1, Migican achieves the state-of-the-art performance across all tasks on MIG-bench, with an average improvement of 21.61% compared to the second-best model, Qwen2-VL-72B (38.88%), despite having significantly fewer parameters.
Note that there is a substantial gap between human performance and that of all MLLMs across all tasks, indicating that MLLMs have significant potential for improvement in free-form MIG. In particular, for 7B-scale models, even advanced multi-image models like InternVL2-8B and Qwen2-VL-7B struggle to perform, particularly in tasks such as multi-view grounding, region locating, and correspondence.
For models equipped with preliminary grounding capabilities, such as mPLUG-Owl3, InternVL2 series, and Qwen2-VL series, their inherent localization ability provides an implicit advantage over other baselines. Furthermore, the proposed single-image CoT method (+CoT) effectively integrates the grounding and multi-image understanding capabilities of the MLLMs where different abilities assist each other in different reasoning steps, achieving comprehensive improvements on multi-image grounding tasks. Moreover, this approach is effective for all the aforementioned models.
@@ 6.3 Results on Multi-Image Understanding Benchmarks
As shown in Table 3, Migician not only establishes its multi-image grounding ability, but also remarkably stimulates its general multi-image understanding ability. In particular, Migician achieves the best average results on the multi-image understanding benchmarks. It surpasses the second-best model (Mantis) on MuirBench by 9.77%, achieving SOTA performance on MMIU and shows a 1.40% improvement on the large-scale MIBench. We attribute this to the training on a mixture of multi-image understanding and grounding data, which indicates that our proposed MGrounding-630k can enhance general multi-image comprehension.
6.4 Results on Single-Image Grounding Bench- marks
@@ 7.1 Effects of Different CoT Strategies
The CoT framework in Section 4.1, after obtaining a referring expression, has the MLLM perform grounding in each image in a polling manner (denoted as single-image CoT), which incurs significant inference overhead. Here, we explore multi-image CoT, where the MLLM directly performs grounding across all images based on the obtained referring expression. As shown in Table 5, multi-image CoT achieves some effectiveness but it still falls significantly behind single-image CoT. In contrast, our proposed Migician is able to perform end-to-end reasoning, offering significant advantages in both efficiency and effectiveness.
7.2 Visual Search in High-Resolution Images
Finding visual details in high-resolution images is a challenging task, and many recent works have explored this area [41]. Typically, these tasks involve images with very high resolution, where the relevant visual information is often quite small, posing significant challenges to the model’s grounding ability. In this section, we analyze and demonstrate that the multi-image grounding capability of Migician can be leveraged to efficiently address this task. Specifically, we slice a single high-resolution image into multiple sub-images and directly transform the problem into a multi-image grounding task. By utilizing the MIG ability of Migician, we can locate the regions relevant to the input question. Afterward, the model combines the identified region with the original image to generate the answer for the input question.
We test this approach on the V*Bench [41] and list the results in Table 3. In the table, we refer to the method that directly asks Migician to answer the question based on the original image as Migicianzero_shot, while Migicianslice denotes the method that transforms this task into a MIG task as mentioned before. Its detailed implementation could be found in Appendix E. The results remarkably demonstrate the effectiveness of using the MIG approach for high-resolution image visual search. Notably, on the Attribute recognition task, Migician even surpasses the specialized visual searching system SEAL [41].

### v3

@@ Abstract
The recent advancement of Multimodal Large Language Models (MLLMs) has significantly improved their fine-grained perception of single images and general comprehension across multiple images. However, existing MLLMs still face challenges in achieving precise grounding in complex multi-image scenarios. To address this, we first explore a Chain-of-Thought (CoT) framework that integrates single-image grounding with multi-image comprehension. While partially effective, it remains unstable and struggles to capture abstract visual information due to its non-end-to-end nature. Therefore, we introduce Migician, the first multi-image grounding model capable of performing free-form and accurate grounding across multiple images. To support this, we present the MGrounding-630k dataset, which comprises data for several multi-image grounding tasks derived from existing datasets, along with newly generated free-form grounding instruction-following data. Furthermore, we propose MIG-Bench, a comprehensive benchmark specifically designed for evaluating multi-image grounding capabilities. Experimental results demonstrate that our model achieves significantly superior multi-image grounding capabilities, outperforming the best existing MLLMs by 24.94% and even surpassing much larger 70B models. Our code, model, dataset, and benchmark are fully open-sourced at https://migician-vg.github.io/.
@@ 1 Introduction
As a result, a question naturally arises: Can we integrate the single-image grounding and multi-image understanding capabilities of existing MLLMs to tackle the MIG task? In this work, we propose a Chain-of-Thought (CoT) framework that first leverages multi-image understanding to generate a textual referring query, and then utilizes it for localization through single-image grounding. This approach has been proven highly effective for MIG tasks, particularly in simple scenarios where textual descriptions are sufficiently distinctive, demonstrating the potential of current MLLMs in handling such tasks.
However, the proposed CoT framework struggles with describing abstract visual semantics in multi-image scenarios, and the two-step process results in a doubling of the inference time. To address this, we further propose Migician, a competitive MLLM capable of free-form and accurate grounding across multiple images, serving as an end-to-end solution for MIG. To progressively establish flexible grounding capabilities, we employ a two-stage training procedure based on our proposed large-scale MIG dataset (MGrounding-630k). First, the grounding ability of Migician is enhanced through a combination of data from MIG tasks and general tasks. Then, Migician is further refined using high-quality free-form MIG instruction data. In addition, to evaluate the challenges of the free-form MIG scenario, we construct a comprehensive multi-image grounding benchmark, MIG-bench, comprising a total of 10 different tasks, 5.9k diverse images and more than 4.2k test instances. We observe a significant gap between the performance of existing mainstream MLLMs and human performance on the MIG-bench. In contrast, Migician can effectively alleviate this gap and improve the performance of free-form MIG.
@@ 4 Methods
In this section, we delve into methods for enabling free-form multi-image grounding capabilities in MLLMs. We begin by exploring a Chain-of-Thought (CoT) framework to elicit the capabilities within existing MLLMs to tackle this task. Then we develop an end-to-end MIG model, Migician, through instruction tuning to overcome the limitations of the CoT framework and achieve enhanced MIG performance.
@@ 4.1 A Chain-of-Thought Framework
As illustrated in Figure 3(a), directly prompting existing MLLMs to perform MIG tasks often leads to significant performance degradation. To better explore the potential of existing models, we design a CoT framework which decouples the MIG task into two stages as shown in Figure 3(b). The model first understands the input images and question, generating a textual referring expression that describes the target object. Next, the model locates the objects in corresponding images using the referring expression from the previous step.
This framework leads to a notable performance improvement on MIG tasks. However, the CoT framework has inherent limitations, such as error propagation due to its multi-step process, which also reduces reasoning efficiency [47]. Additionally, as illustrated in Figure 5(c), many scenarios require grounding through abstract visual semantics across multiple images, which cannot be effectively captured through textual expressions. More failure types are detailed in Appendix C. This highlights the need for an end-to-end model capable of directly performing the MIG task.
@@ 4.2 Data Construction
By analyzing the tasks and annotation types of existing datasets, we identify multiple multi-image grounding (MIG) tasks whose data could be derived through transformation of the existing. Specifically, we collect and organize data from existing sources, combining or automatically synthesizing single-image annotations to create datasets for 6 types of MIG tasks. Each task contains over 70k examples, resulting in a total of 530k training samples. The details of these task data are in Appendix D.1.
Synthesizing Free-form MIG Data.
The data obtained through the aforementioned methods still do not fully meet the requirements for free-form MIG. To acquire MIG data with richer and more diverse formats, which would enhancing the model’s instruction-following and flexible grounding capabilities, we design a MIG data synthesis pipeline. This pipeline uses the Object365 [38] images with object annotations, select multiple images as a group, and generate high-quality instructions for multi-image grounding. Specifically, we first employ Qwen2-VL-72B [41] to generate captions of each individual image and then perform error filtering and refinement on the annotated bounding boxes. Next, we prompt Qwen2.5-72B [46] to automatically generate high-quality, free-form MIG question-answering pairs by integrating information from multiple images. To optimize the selection of appropriate image groups, we adopt different image grouping methods, including random selection, selection of images with common objects, and grouping images based on CLIP similarity to select semantically similar images for each. Using these methods, we generate a total of 100k Free-Form MIG data. Detailed information can be found in Appendix D.2.
4.3 Instruction Tuning for MIG
Two-Stage Training.
To effectively equip the model with free-form MIG capabilities, we propose a two-stage training approach. In the first stage, the model learns to perform multi-image grounding by training on the six representative MIG tasks of MGrounding-630k, acquiring the ability to simultaneously comprehend multiple images and execute visual grounding. In the second stage, the model is further fine-tuned on free-form MIG instruction data in MGrounding-630k, enabling it to adapt to more flexible and diverse instruction types and transfer the MIG skills learned in the first stage to a broader range of scenarios. To prevent the model from forgetting its existing capabilities during training, we also incorporate single-image understanding, multi-image understanding, and single-image grounding data into each training stage. For more details please refer to the Appendix E.
@@ 4.3 Instruction Tuning for MIG
After the second stage of fine-tuning, we observe a trade-off between model performance and flexibility: while the model adapts to the free-form MIG instructions, there is a performance drop in common multi-image grounding tasks. To better balance these two aspects, we adopt the model merging technique [15], averaging the model weights obtained from both training stages as the final weights. We find this approach mitigates the performance loss in common MIG tasks while preserving the ability to follow free-form MIG instructions effectively.
5 MIG-Bench
We introduce MIG-Bench, a manually curated benchmark designed to evaluate the MIG ability of current MLLMs. It comprises 5.9k images, and 4.3k testing instances, covering 10 distinct tasks shown in Figure 2 with details in Appendix A. The distribution of these tasks is illustrated in Figure 4.
MIG-Bench is manually constructed from multiple data sources. Initially, we select annotated data examples from existing datasets and adapt them for the MIG task. We collect challenging examples from Objects365 [38] for Common Object Grounding and Region Locating. We use examples that exhibits significant movement from GOT-10k [14] for Object Tracking. For Multi-View Grounding, we utilize the rich annotations from Ego4D [12]. The Static Difference task is sourced from MagicBrush [53]. We combine multiple examples from GranD [36] to form a group for Group Grounding. Additionally, for tasks such as Reasoning Grounding, Correspondence, Referring Grounding, and Robust Diff Grounding, which lack suitable existing datasets, we collect both web images and manually captured photos, and annotate them with well-educated annotators who are thoroughly trained and fully understand these tasks.
All instances are reviewed by two different human annotators to guarantee the quality of MIG-Bench. This includes removing instances with incorrect annotations, ensuring that the questions are answerable, filtering out overly simplistic questions, and refining ambiguous queries. We also invite five volunteers to answer the questions to evaluate human performance on this benchmark (detailed in Appendix F). As shown in Table 1, the average accuracy of human responses is 97.18%, indicating that the task is easy for humans and further demonstrating the high quality of MIG-Bench.
Unlike existing benchmarks, MIG-Bench introduces the grounding task in a multi-image setting, thereby addressing the gap in previous benchmarks that are unable to measure the free-form MIG capabilities of MLLMs. A detailed comparison with other benchmarks is provided in Appendix B.
6 Experiments
6.1 Implementation Details
Migician undergoes development based on the Qwen2-VL-7B [41] foundation model with a global batch size of 48, a total of 25,000 steps for the two-stage training procedure, and a learning rate of 5e-6, using 8×A100-80G GPUs. For the evaluation in our proposed MIG-Bench, we use the conventional metric $\text{Acc}_{0.5}$ in referring expression comprehension [20]. This metric measures the accuracy of object localization, defining a prediction as correct if the Intersection over Union (IoU) with the ground truth is greater than 0.5.
6.2 Results on MIG-Bench
As shown in Table 1, Migican achieves the state-of-the-art performance across all tasks on MIG-bench, with an average improvement of 24.94% compared to the second-best model, Qwen2-VL-72B (38.88%), despite having significantly fewer parameters. Note that there is a substantial gap between human performance and that of all MLLMs across all tasks, indicating that MLLMs have significant potential for improvement in free-form MIG. In particular, for 7B-scale models, even advanced multi-image models like InternVL2-8B and Qwen2-VL-7B struggle to perform, particularly in tasks such as multi-view grounding, region locating, and correspondence.
For models equipped with grounding capabilities, such as mPLUG-Owl3, InternVL2 series, and Qwen2-VL series, they demonstrate an advantage over other baselines. Furthermore, the proposed single-image CoT method (+CoT) effectively integrates the grounding and multi-image understanding capabilities of the MLLMs where different abilities assist each other in different reasoning steps, achieving comprehensive improvements on multi-image grounding tasks. Moreover, this approach proves effective for all the aforementioned models.
@@ 6.2 Results on MIG-Bench
Model MME MMBench V* Bench
Closed-Source Model
GPT-4V 1926.6 81.0 54.97
Gemini-Pro 2148.9 73.6 48.16
Claude-3.5 1920.0 79.7
Open-Source Model
LLaVA-1.5 1510.7 64.3 48.68
InternVL2-8B 2210.3 81.7 43.07
MiniCPM-V 2.6 2024.6 77.2 52.67
SEAL 1128.9 33.1 75.39
LLaVA-OV-7B 1998.0 80.9
Mantis 1806.4 75.7
Migician 2244.7 80.0 72.30
Table 3: The performance of models on various single-image benchmarks, where Migician consistently exhibits strong capabilities.
@@
Setting Multi-image General Benchmarks MIG
MuirBench BLINK MIBench Mantis MMIU
@@
Table 6: The ablation study on the contribution of different data subsets.
Model Easy Medium Hard
# Instances 2471 1430 395
InternVL2-8B 44.69 13.92 1.77
Qwen2-VL-7B 30.31 22.24 0.00
InternVL2-8B+CoT 67.10 7.06 0.25
Qwen2-VL-7B+CoT 71.02 10.70 0.76
Migician 76.00 52.10 29.37
Table 7: Performance comparison across varying difficulty levels on MIG-Bench.
@@ 6.3 Results on Multi-Image Understanding Benchmarks
As shown in Table 3, Migician not only establishes its multi-image grounding ability, but also remarkably stimulates its general multi-image understanding ability. In particular, Migician achieves the best average results on the multi-image understanding benchmarks. It surpasses the second-best model (Mantis) on MuirBench by 9.19%, achieving SOTA performance on both MMIU and MIBench. We attribute this to the training on a mixture of multi-image understanding and grounding data, which indicates that our proposed MGrounding-630k dataset can effectively enhance general multi-image comprehension.
6.4 Results on Single-Image Benchmarks
Table 3 lists the empirical results on typical single-image understanding benchmarks including MME [10] and MMBench [29], suggesting that Migician retains strong single-image understanding capacities. Specifically, on the MME benchmark, Migician surpasses notable MLLMs like InternVL2-8B and MiniCPM, while exhibiting equally strong performance with close-source models on MMBench. Notably, contrasting with specialized multi-image models such as LLaVA-OV and Mantis, whose single image ability has largely degenerated, Migician poses significant advantage over them, achieving comprehensive capability maintenance. Our model also maintains strong performance on single-image REC tasks, as shown in Table 5.
Furthermore, we find that the MIG ability of Migician can be leveraged to address the task of finding visual details in high-resolution images, such as V*Bench [44]. Specifically, we split a single high-resolution image in V*Bench into multiple sub-images and transform the problem into a MIG task (detailed in Appendix F). Results show that Migician can generalize well to this out-of-distribution setting, performing on par with the specialized visual searching system SEAL [44].
@@ 7.1 Effects of Different CoT Strategies
The CoT framework in Section 4.1, after obtaining a referring expression, has the MLLM perform grounding in each image in a polling manner (denoted as single-image CoT), which incurs significant inference overhead. Here, we explore multi-image CoT, where the MLLM directly performs grounding across all images based on the obtained referring expression. As shown in Table 5, multi-image CoT achieves some effectiveness but it still falls significantly behind single-image CoT. In contrast, our proposed Migician is able to perform end-to-end reasoning under the multi-image setting, offering significant advantages in both efficiency and effectiveness.